if a toy rocket is launched vertically upward from ground level with an initial velocity of 128 feet per second, then its height h after t seconds is given by the equation h(t)= -16t[square] +128t (if air resistence is neglected). How long will it take for the rocket to return to the ground? After how many seconds will the rocket be 112 feet above the ground? how long will it take the rocket to hit its maximum height? What is the maximum height?
Answer by lwsshak3(11628)

if a toy rocket is launched vertically upward from ground level with an initial velocity of 128 feet per second, then its height h after t seconds is given by the equation h(t)= -16t[square] +128t (if air resistance is neglected). How long will it take for the rocket to return to the ground? After how many seconds will the rocket be 112 feet above the ground? how long will it take the rocket to hit its maximum height? What is the maximum height?
***
h(t)= -16t^2 +128t
returning to ground:
0=-16t(t-8)
t=0
or
t=8
How long will it take for the rocket to return to the ground? 8 sec
..
112 ft above ground:
112=-16t^2+128t
-16t^2+128t-112=0
-t^2+8t-7=0
(-t+1)(t-7)=0
t=1
or
t=7
After how many seconds will the rocket be 112 feet above the ground? after 1 sec on the way up and after 7 sec on the way down
..
maximum height
h(t)= -16t^2 +128t
complete the square:
h(t)= -16(t^2 -8t+16)+256
h(t)=-16(t-4)^2+256
This is an equation of a parabola that opens downwards with vertex at (4,256)
Rocket will reach its maximum height of 256 ft 4 sec after launch